# Continual Unsupervised Out-of-Distribution Detection

### v1

@@ 3.1.2 Few-shot learner $s^{-}$:
where the factor $\lambda$ controls the influence of $\sigma^{\textrm{out}}$ in the final score. The value of $\lambda$ depends on the confidence levels of the detectors, which, in turn, rely on the contents in $\mathcal{D}^{\textrm{in}}$ and $\mathcal{D}^{\textrm{out}}$. For instance, a small number of samples in $\mathcal{D}^{\textrm{out}}$ will result in high uncertainty in the estimations of $\sigma^{\textrm{out}}$, which should be compensated with a small $\lambda$, and vice versa.
@@
When working with image data, we proceed layer by layer, as previously discussed for the OOD detector. In particular, we build a few-shot learner $s^{-}_{\ell}$ per each layer $\ell$ of the feature extractor $\mathbf{f}$, and the final score is the sum of the layer-wise scores,
@@ 4 Experiments
In the following section, we describe our experimental setup to evaluate continual U-OOD. To assess the performance of our approach, seven baselines are evaluated on three proposed benchmarks. In addition, we also perform an ablation study to identify contributing aspects of our method and compare our setting to outlier exposure, i.e., having access to a large dataset of auxiliary outliers from the unconstrained $\mathcal{P}_{\textrm{out}}$.
@@ 4.1.1 Baselines:
We compare our method to seven baselines that comprise the top-performing methods from related fields: AdaODD [54], ETLT [10], and SAL [8] from OOD test-time adaptation, BCE from outlier exposure [30], the Mahalanobis difference (MDiff) [47] from few-shot OOD, and HSC [30] from semi-supervised OOD. Furthermore, we include MahaAD [41] as a U-OOD baseline. All methods use the same ImageNet pre-trained backbone.
To ensure fairness, our method and all baselines use a grid search procedure to set hyperparameters. We measure the performance of each method on the CIFAR10 experiment Plane:Rest and select the highest-performing configuration over the grid. The ranges considered for each hyperparameter in the grid search are given in the supplementary material.
4.1.2 Datasets:
We introduce three continual U-OOD benchmarks on which we compare our method against the baselines under varying settings:
  1. 1.
    Multi-class: we adapt two common benchmarks with CIFAR10 and CIFAR100 [26] as the in-distribution [8, 23]. For both datasets, we run experiments with SVHN [33], resized LSUN [53], resized TinyImageNet [29], Places365 [55], and Describable Textures [4] as OOD. For LSUN and TinyImageNet, we use the fixed versions from [50]. Each experiment has $T=5$ steps with $K=100$ unlabeled samples and contamination $\pi=0.1$.
  2. 2.
    One-class: continual versions of the ten quintessential one-class experiments on CIFAR10 [26], where one of the classes is the in-distribution and the other nine comprise the out-distribution [7, 38, 50]. Each experiment has $T=10$, $K=40$ and $\pi=0.25$.
  3. 3.
    High-resolution: three experiments with high-resolution images selected from [7]. The first trains on healthy chest X-rays and uses pathological chest X-rays as OOD [51]. The second is Real:Infograph on DomainNet [37], which uses real images as in and infographs as OOD. The third has healthy retinal fundus photographs as the in-distribution, with the strongest level of diabetic retinopathy as OOD [13]. For all three cases, we set $T=5$, $K=50$, and $\pi=0.2$.
There is no overlap between the test samples utilized for evaluating the quality of the methods and those belonging to $\mathcal{D}_{t}^{\textrm{deploy}}$.
Table 1: Comparative evaluation on Multi-class. We report the mean of the AUF ($\downarrow$) and AUA ($\uparrow$) over five trials. Bold and underlined indicate best and second best, respectively. On average, our method obtains the best performance.
SVHN LSUN ImgNet Places DTD Mean
AUF AUA AUF AUA AUF AUA AUF AUA AUF AUA AUF AUA
CIFAR10
SAL 91.4 63.3 93.2 52.6 91.6 55.6 82.7 62.1 92.2 53.6 89.9 56.0
HSC 94.6 53.4 86.4 61.8 80.0 64.8 33.2 90.5 87.5 60.7 76.3 66.2
ETLT 87.8 56.7 85.4 66.0 68.2 71.9 27.9 88.9 85.1 58.2 70.9 68.3
MahaAD 63.0 86.2 84.5 71.7 64.6 76.7 22.4 92.4 85.0 64.3 63.9 78.3
BCE 6.0 98.3 63.1 76.5 77.8 68.3 26.2 92.0 83.1 64.0 51.2 79.8
AdaODD 67.4 83.7 68.2 82.1 47.1 86.0 15.8 95.7 70.7 79.0 53.8 85.3
MDiff 1.3 99.7 58.2 85.7 61.3 78.8 13.7 96.1 75.0 72.9 41.9 86.6
ours 3.0 99.2 51.4 88.9 53.4 83.7 11.1 97.2 65.7 81.4 36.9 90.1
CIFAR100
SAL 93.5 54.9 84.8 61.4 94.4 52.2 88.0 56.0 92.8 54.8 90.7 55.9
HSC 99.8 34.8 95.1 53.1 91.8 53.0 56.8 81.5 96.0 50.4 87.9 54.6
ETLT 98.8 35.6 98.2 50.0 90.7 58.0 47.3 79.8 96.6 44.8 86.3 53.6
MahaAD 96.8 62.1 97.0 56.0 85.3 63.5 40.2 84.3 95.8 49.7 83.0 63.1
BCE 33.2 86.1 85.7 63.9 92.2 53.9 56.2 77.8 92.6 56.2 72.0 67.6
AdaODD 93.8 55.7 93.0 61.6 78.0 73.4 37.3 87.5 90.5 63.0 78.5 68.2
MDiff 25.6 95.3 88.9 72.1 83.7 67.5 32.5 89.0 93.0 60.9 64.7 77.0
ours 55.8 85.3 88.4 74.4 79.9 70.7 28.3 91.6 89.7 66.6 68.4 77.7
@@ 4.1.3 Evaluation metrics:
The two metrics commonly used to measure the performance of OOD methods are the false positive rate at 95% true positive rate (FPR@95) and the area under the receiver operating characteristic curve (AUC) [9, 29, 49]. In contrast with previous works, we are interested in measuring the quality of the OOD detector over time. As such, we propose two metrics that consider this aspect: the Area Under the FPR@95 curve (AUF) and the Area Under the AUC curve (AUA). These are computed by first evaluating the FPR@95 and AUC at every timestep and plotting the resulting FPR@95/AUC curves with respect to $t$. Then, the AUF and AUA are given by the area under the FPR@95 and AUC curves, respectively, normalized by the time elapsed.
@@ 4.1.4 Implementation details:
Our pre-trained model $\mathbf{f}$ is a ResNet-18 [15] pre-trained on ImageNet [5]. For $s^{-}$, we extract features from $L=4$ layers at the end of every ResNet block and normalize the features of the final layer following [10, 39]. The binary classifier $s^{+}$ also uses a ResNet-18 architecture. It is trained with Adam [24] using a learning rate of $10^{-5}$ and a batch size of 256 for ten epochs on CIFAR10 and an equivalent number of iterations on the other datasets. We apply data augmentation in the form of random resized crops, color jitter, and horizontal flips and initialize the binary classifier with the weights from the previous timestep. At every step, the elements of $\mathcal{D}_{t}^{\textrm{deploy}}$ are re-labeled with the current model $s_{t}(\mathbf{x})$ using a threshold. The threshold is determined on the training data such that 95% is considered in-distribution. We standardize the scores of $s^{+}$ and $s^{-}$ before combining them to ensure they have similar scales.
Table 2: Comparative evaluation on One-class. We report the mean of the AUF ($\downarrow$) over five trials (AUA in supplementary). Bold and underlined indicate best and second best, respectively. Our method is the best overall.
Plane Car Bird Cat Deer Dog Frog Horse Ship Truck Mean
SAL 82.9 46.6 85.2 80.0 78.5 78.8 64.7 78.4 63.5 40.4 69.9
HSC 64.2 57.1 86.6 81.6 71.9 79.3 53.3 61.2 48.9 52.1 65.6
ETLT 47.8 23.1 61.7 69.3 34.8 48.9 40.7 42.4 27.2 20.3 41.6
MahaAD 49.6 27.8 65.0 74.9 36.4 51.5 43.9 46.4 33.0 20.6 44.9
BCE 41.6 29 53.8 54 47.5 52 33.2 50 31 29.1 42.1
AdaODD 35.7 13.0 56.5 59.1 26.9 37.8 25.1 22.0 21.6 10.8 30.9
MDiff 28.3 18.6 45.7 50.3 31 37.2 25.9 31.6 19.4 15.6 30.4
ours 29.5 14.0 43.2 49.9 26.4 34.0 22.6 22.0 17.8 12.5 27.2
Table 3: Comparative evaluation on High-resolution. We report the mean of the AUF ($\downarrow$) and AUA ($\uparrow$) over five trials. Bold and underlined indicate best and second best, respectively. We obtain the best or second performance on all datasets.
NIH DN DRD
AUF AUA AUF AUA AUF AUA
SAL 89.5 56.8 90.5 54.9 66.0 74.4
HSC 66.0 79.7 79.6 70.2 80.7 71.9
ETLT 56.1 81.6 85.8 70.6 62.9 78.7
MahaAD 53.8 85.3 83.4 73.9 68.7 78.5
BCE 68.8 67.9 45.5 80.5 43.2 86.5
AdaODD 46.8 87.0 89.8 76.3 56.2 83.6
MDiff 49.7 82.4 28.6 92.2 55.0 84.6
ours 39.3 89.8 36.3 93.5 53.9 85.1
@@ 4.2 Results
Tab. 1 reports the results for Multi-class. Out of all methods, SAL, HSC, and ETLT are unable to outperform the unsupervised MahaAD baseline in any scenario. Unsurprisingly, we find the gradient-based filtering in SAL unable to detect OOD samples with the ImageNet pre-trained network. We also observe that the supervised classifier BCE is generally inconsistent; for instance, it achieves excellent results when SVHN is the OOD distribution but performs below average in the other cases. In contrast, AdaODD scores consistently high in all experiments except for SVHN but is, in turn, outclassed by MDiff. More broadly, our method is consistently among the top performers, reaching the best score for LSUN, Places, and DTD with both CIFAR10 and CIFAR100 as the in-distribution and scoring second-best in three out of the four remaining experiments.
In the One-class benchmark (Tab. 2), HSC also struggles, while ETLT consistently outperforms MahaAD, unlike on the Multi-class benchmark. AdaODD and MDiff again perform well, but which one is better between them depends on the experiment. Here, we also find our method to outperform all cases but three, where it performs second-best. On High-resolution, Tab. 3 reports similar trends for our method ranking best or second in all six cases. Interestingly, BCE, which performs poorly on average across all three benchmarks, reaches the best performance on DRD.
The performance evolution for some of the best methods on the Bird:Rest and Healthy:Pathological experiments are shown in Fig. 3. All methods benefit from incorporating unlabeled samples, with the exception of AdaODD, whose optimal hyperparameters from the grid search assign low importance to the unlabeled data.
@@
Refer to caption Refer to caption
(c) (d)
Figure 3: Performance curves for top-performing methods. We show the AUC and FPR@95 curves for Bird:Rest in (a) and (b), and for Healthy:Pathological in (c) and (d). Our method achieves the best results and can improve results with a few unlabeled test samples.
4.3 Ablations
We ablate our design choices by showing that (1) our OOD scoring function combining the Mahalanobis distance with k-nearest neighbors (MkNN) outperforms both kNN and MahaAD for U-OOD detection and (2) our few-shot learner $s^{-}$ outperforms unscaled scoring functions on few-shot OOD. To do so, we run experiments with our method on the standard one-class CIFAR10 benchmark. For few-shot OOD, we run experiments in the 5- and 10-shot settings. Here, $n$-shot refers to using $n$ ground-truth OOD samples from the test set. For CIFAR10, this implies that the 5-shot setting will, by definition, not include an example of all OOD classes, as there are nine classes in the OOD distribution. Similarly, the 10-shot case will often not cover the whole class distribution due to the random selection of the 10 OOD samples.
@@ 4.3 Ablations
We show the results for U-OOD in Tab. 4. MkNN outperforms the kNN scoring by 5.9 AUC and MahaAD by 1.6 AUC, showcasing its practical usefulness as MahaAD is one of the top-performing U-OOD detectors [7]. From Tab. 5, we find that equipping MDiff with our confidence scaling, which we label as ($s^{-}_{Maha}$), improves the results by 6.5 AUC and 2 AUC in the 5- and 10-shot case, respectively. Furthermore, we improve the results by a further 0.6 and 0.8 AUC by using MkNN.
Table 4: Comparing U-OOD performance on CIFAR10. We report the AUC over one run as the methods are deterministic. Bold indicates the best method. MkNN is the best scoring function in all ten U-OOD experiments.
Plane Car Bird Cat Deer Dog Frog Horse Ship Truck Average
kNN 80.0 75.5 86.3 79.6 82.7 82.4 87.7 76.5 82.4 84.3 81.7
MahaAD 86.2 78.0 89.3 88.7 85.6 80.8 92.1 79.5 90.0 90.0 86.0
MkNN (ours) 88.3 80.3 89.9 89.6 87.3 84.2 92.7 81.9 90.9 90.3 87.6
Table 5: Comparing few-shot OOD performance on CIFAR10. We report the mean AUC over five runs. Bold indicates the best method. Our few-shot detector $s^{-}_{MkNN}$ is best overall.
Plane Car Bird Cat Deer Dog Frog Horse Ship Truck Average
5-shot
kNN 85.8 79.8 77.5 70.4 74.6 72.0 80.5 64.6 80.4 70.7 75.6
MDiff 90.2 92.1 80.0 75.5 83.0 82.6 89.7 76.1 88.9 89.1 84.7
$s^{-}_{Maha}$ (ours) 91.7 94.4 84.7 83.3 92.2 90.5 92.4 92.4 94.1 96.4 91.2
$s^{-}_{MkNN}$ (ours) 93.1 95.2 87.0 84.3 92.4 90.7 92.6 92.6 93.5 96.3 91.8
10-shot
kNN 87.6 82.0 79.5 74.5 79.0 77.0 85.7 70.4 83.2 78.9 79.8
MDiff 92.8 94.5 84.1 80.7 90.6 86.6 93.9 88.4 93.3 93.9 89.9
$s^{-}_{Maha}$ (ours) 92.5 94.8 85.7 84.2 92.8 90.9 93.5 92.5 95.2 96.7 91.9
$s^{-}_{MkNN}$ (ours) 93.9 95.7 87.4 85.0 93.7 91.6 94.1 93.2 95.2 96.9 92.7
4.4 Continual U-OOD vs. Outlier Exposure
To illustrate the importance of continual U-OOD, we show how using a small number of unlabeled samples obtained iteratively during deployment is beneficial compared to the standard OE setting, which uses a large dataset of auxiliary outliers. For OE, we take 50’000 images from TinyImageNet as the outliers and train a binary classifier between the in-distribution and out-distribution [30]. We compare this to the results obtained with our method using the continual U-OOD setting of the previous experiments.
Fig. 4 shows that even when pseudo-labeling $\mathcal{D}_{t}^{\textrm{deploy}}$ containing only 10 OOD samples, continual U-OOD is already far more effective than OE with 50’000 samples, with the gap widening as $\left|\mathcal{D}_{t}^{\textrm{deploy}}\right|$ increases. This is especially true when there is a mismatch between OE and OOD distributions, such as for the medical images in Healthy:Pathological (Fig. 4b).
Refer to caption Refer to caption
(a) (b)
Figure 4: Comparing continual U-OOD with Outlier Exposure. We show the AUA ($\uparrow$) for Bird:Rest in (a) and Healthy:Pathological in (b). Continual U-OOD is far more effective than OE, especially in the case of a mismatch between OE and OOD distributions.
4.5 Varying Contamination Ratios
We investigate the effectiveness of our proposed method compared to two baselines under varying contamination ratios $\pi$ while keeping the number of OOD samples per step fixed at ten. We show results on CIFAR10:TinyImageNet and CIFAR10:DTD in Fig. 5. Our method achieves a better AUA for all contamination levels. As expected, the methods benefit from having a higher fraction of unlabeled OOD samples in the test set. Nonetheless, with a contamination ratio as low as 0.05 (i.e., $\mathcal{D}_{t}^{\textrm{deploy}}$ contains 200 unlabeled samples, of which only ten are OOD), we still outperform the unsupervised baseline by 2.5 AUA on CIFAR10:TinyImageNet and 9.1 AUA on CIFAR10:DTD.
Refer to caption Refer to caption
(a) (b)
Figure 5: Comparing methods under varying contamination ratios. We show the AUA per contamination ratio for CIFAR10:TinyImageNet in (a) and CIFAR10:DTD in (b). Our method outperforms the baselines at all contamination levels.

### v2

@@ 2.1.2 Few-shot learner $s^{-}$:
where the factor $\lambda$ controls the influence of $\sigma^{\textrm{out}}$ in the final score. The value of $\lambda$ depends on the confidence levels of the detectors, which, in turn, rely on the contents in $\mathcal{D}^{\textrm{in}}$ and $\mathcal{D}^{\textrm{out}}$.
@@
When working with image data, we proceed layer by layer, as previously discussed for the U-OOD detector. In particular, we build a few-shot learner $s^{-}_{\ell}$ per each layer $\ell$ of the feature extractor $\mathbf{f}$, and the final score is the sum of the layer-wise scores,
@@ 3.1 Experimental set-up
3.1.1 Datasets:
We introduce three IDE benchmarks to compare methods comprising various modalities, contamination ratios, and other settings. The samples for evaluating the methods and those in $\mathcal{D}_{t}^{\textrm{deploy}}$ do not overlap.
  1. 1.
    NIH [26]: Training: 4’261 healthy chest X-rays. Testing: 250 healthy scans as ID and 250 pathological chest X-rays as OOD. All methods see $T=10$ steps of $K=50$ test samples, with contamination $\pi=0.2$.
  2. 2.
    MURA [16]: Training: 5’106 musculoskeletal radiographs of fingers. Testing: 250 finger scans as ID with scans of elbows, forearms, hands, humeri, shoulders, and wrists as OOD, where $T=5$, $K=100$, and $\pi=0.1$.
  3. 3.
    DRD [6]: Training: 25’809 healthy retinal fundus photographs as ID. Testing: 250 healthy scans as ID, with 250 scans of strongest level of diabetic retinopathy as OOD, where $T=5$, $K=50$, and $\pi=0.1$.
@@ 3.1.2 Baselines:
We compare our method to seven baselines that comprise the top-performing methods from related fields: AdaODD [28], ETLT [4], SAL [3], BCE [13], the Mahalanobis difference (MDiff) [22], and HSC [13]. Furthermore, we include MahaAD [19] as a U-OOD baseline. All methods use the same ImageNet pre-trained backbone. To ensure fairness, all methods use the same grid search procedure to set their hyperparameters. We measure the performance of each method on the CIFAR10 experiment Plane:Rest and select the highest-performing configuration over the grid.
@@ 3.1.3 Evaluation metrics:
In contrast with previous works, we are interested in measuring the quality of the OOD detector over time. As such, we propose two metrics that consider this aspect: the Area Under the FPR@95 curve (AUF) and the Area Under the AUC curve (AUA). These are computed by first evaluating the FPR@95 and AUC at every timestep and plotting the resulting FPR@95/AUC curves with respect to $t$. Then, the AUF and AUA are given by the area under the FPR@95 and AUC curves, respectively, normalized by the time elapsed.
@@ 3.1.4 Implementation details:
For $s^{-}$, we extract features from $L=4$ layers at the end of every ResNet-18 block and normalize the features of the final layer following [4, 18]. The binary classifier $s^{+}$ also uses a ResNet-18 architecture. It is trained with Adam [10] using a learning rate of $10^{-5}$ and batch size 256 for ten epochs on NIH and an equivalent number of iterations on the other datasets. We apply data augmentation in the form of random resized crops, color jitter, and horizontal flips and initialize the binary classifier with the weights from the previous timestep. At every step, the elements of $\mathcal{D}_{t}^{\textrm{deploy}}$ are re-labeled with the current model $s_{t}(\mathbf{x})$ using a threshold. The threshold is determined on the training data such that 95% is considered in-distribution. We standardize the scores of $s^{+}$ and $s^{-}$ before combining them to ensure they have similar scales. The hyperparameters $\beta$ and $\gamma$ were set via the grid search to $1/300$ and $3$, respectively. We found CSO robust to their settings, as shown in the next section.
Table 1: Comparative evaluation. We report the mean of the AUF ($\downarrow$) and AUA ($\uparrow$) over five trials. Bold and underlined indicate best and second best, respectively. Our method obtains the best performance overall.
NIH MURA DRD Mean
AUF AUA AUF AUA AUF AUA AUF AUA
SAL 91.5$\pm$6.3 58.3$\pm$12.1 90.0$\pm$8.3 56.6$\pm$9.2 77.4$\pm$8.9 67.3$\pm$7.9 86.3 60.7
HSC 70.4$\pm$0.9 78.6$\pm$0.2 68.8$\pm$6.5 79.8$\pm$2.3 81.8$\pm$2.0 71.6$\pm$1.3 73.7 76.7
ETLT 54.9$\pm$4.0 82.0$\pm$1.0 61.5$\pm$4.9 86.5$\pm$0.7 64.8$\pm$2.0 78.3$\pm$1.1 60.4 82.3
BCE 60.1$\pm$16.7 74.0$\pm$9.6 56.8$\pm$8.2 84.3$\pm$3.1 60.5$\pm$16.9 80.3$\pm$7.3 59.1 79.5
MahaAD 53.8$\pm$3.9 85.3$\pm$0.9 47.0$\pm$4.5 90.8$\pm$0.7 71.4$\pm$3.9 78.1$\pm$1.0 57.4 84.7
AdaODD 46.9$\pm$5.5 87.0$\pm$0.9 48.3$\pm$7.9 88.6$\pm$1.3 57.8$\pm$4.4 83.5$\pm$1.1 51.0 86.4
MDiff 48.1$\pm$4.4 83.0$\pm$2.1 42.9$\pm$6.6 91.3$\pm$1.1 60.3$\pm$4.6 82.7$\pm$3.1 50.4 85.7
CSO (ours) 34.9$\pm$6.1 90.9$\pm$1.7 37.5$\pm$5.6 91.8$\pm$0.9 61.9$\pm$5.6 83.1$\pm$1.6 44.8 88.6
@@ 3.2 Results
Tab. 1 reports the results on our benchmark. Of all methods, SAL, HSC, ETLT, and BCE do not outperform the unsupervised MahaAD baseline on average. BCE is especially inconsistent; for instance, it achieves excellent results on DRD but performs below average in other cases, demonstrating the need for a more robust method. In contrast, AdaODD and MDiff score consistently high in all experiments. Nonetheless, they are outclassed by CSO, which reaches the best score for NIH and MURA and the best overall performance by 5.6 AUF and 2.2 AUA compared to the next-best method.
The performance evolution over time for some of the best methods on NIH is shown in Fig. 2(a). All methods benefit from incorporating unlabeled samples, with our method improving fastest. The only exception is AdaODD, whose optimal hyperparameters from the grid search assign low importance to the unlabeled NIH data.
3.3 Ablations
3.3.1 Hyperparameter sensitivity
From the grid search conducted on Plane:Rest, we find CSO to be robust to its main hyperparameters $\beta$ and $\gamma$: all combinations with $\gamma\in[1,3,5]$ and $\beta\in[1/100,1/300,1/500]$ achieve between $92.6$ and $93.2$ AUA.
3.3.2 Contamination ratio
We probe the effectiveness of CSO compared to the three best baselines under varying contamination ratios $\pi$ while keeping the number of OOD samples per step fixed at ten in Fig. 2(b). As expected, all iterative methods benefit from having a higher fraction of unlabeled OOD samples in the test set. Nonetheless, CSO achieves the best AUF for all contamination levels.
@@ 3.3.2 Contamination ratio
Figure 2: Ablation curves. In (a), we show how the FPR@95 evolves over time for the best methods on NIH. Our method achieves the best results, already after one iteration. In (b), we compare methods by AUF under varying contamination ratios on NIH. Our method consistently outperforms the baselines at different contamination levels.
3.3.3 Scoring function
We ablate our design choices by showing that (1) MkNN outperforms both kNN and MahaAD for U-OOD detection and (2) $s^{-}$ outperforms unscaled scoring functions on few-shot OOD. To do so, we run experiments with our method on the standard one-class CIFAR10 benchmark.
@@ 3.3.3 Scoring function
From Tab. 2(left), MkNN outperforms the kNN scoring by 5.9 AUC and MahaAD by 1.6 AUC, showcasing its practical usefulness over top-performing U-OOD detectors [2]. Tab. 2(right) shows that equipping MDiff with our confidence scaling, which we label as $s^{-}_{Maha}$, already improves few-shot results. We improve the results by a further 0.6 and 0.8 AUC using MkNN. These results on natural images also confirm the usefulness of our method beyond medical settings.
Table 2: Ablation study on CIFAR10. (left) U-OOD performance in AUC over one run as the methods are deterministic. MkNN outperforms the other U-OOD scoring functions. (right) Comparing few-shot OOD performance. We report the mean AUC over five runs. $n$-shot refers to using $n$ ground-truth OOD samples. The confidence scaling is important to achieve the best results.
AUC
kNN 81.7
MahaAD 86.0
MkNN (ours) 87.6
5-shot AUC 10-shot AUC
kNN 75.6 79.8
MDiff 84.7 89.9
$s^{-}_{Maha}$ (ours) 91.2 91.9
$s^{-}_{MkNN}$ (ours) 91.8 92.7